### v3

# FinTruthQA: A Benchmark Dataset for Evaluating the Quality of Financial Information Disclosure

Abstract
Accurate and transparent financial information disclosure is essential in accounting and finance, fostering trust and enabling informed investment decisions that drive economic development. Among many information disclosure platforms, the Chinese stock exchanges’ investor interactive platform provides a novel and interactive way for listed firms to disclose information of interest to investors through an online question-and-answer (Q&A) format. However, it is common for listed firms to respond to questions with limited or no substantive information, and automatically evaluating the quality of financial information disclosure on large amounts of Q&A pairs is challenging. In this study, our interdisciplinary team of AI and finance professionals proposed FinTruthQA, a benchmark designed to evaluate advanced natural language processing (NLP) techniques for the automatic quality assessment of information disclosure in financial Q&A data. It comprises 6,000 real-world financial Q&A entries and each Q&A was manually annotated based on four key evaluation criteria: question identification, question relevance, answer readability, and answer relevance. We benchmarked various NLP techniques on FinTruthQA, including statistical machine learning models, pre-trained language model and their fine-tuned versions, as well as large language models (LLMs). Experiments showed that existing NLP models have strong predictive ability for question identification and question relevance tasks, but are suboptimal for answer readability and answer relevance tasks. By establishing this benchmark, we provide a robust foundation for the automatic evaluation of information disclosure, demonstrating how AI can be leveraged for social good by promoting transparency, fairness, and investor protection in financial disclosure practices. FinTruthQA can be used by auditors, regulators, and financial analysts for real-time monitoring and data-driven decision-making, as well as by researchers for advanced studies in accounting and finance, ultimately fostering greater trust and efficiency in the financial markets.
1 Introduction
Within the financial market, improving transparency and timeliness in information disclosure can mitigate information friction, empower investors to make informed decisions, and promote market efficiency. According to Lee and Zhong Lee and Zhong (2022), retail investors struggle to integrate the existing public information, and increased activities on the online platform are associated with more active market transactions, such as greater trading volume, return volatility, and market liquidity. New interactive investor platforms (such as online forums established by the Shanghai Stock Exchange (SSE) and Shenzhen Stock Exchange (SZSE)) offer a unique avenue for retail investors to engage directly with listed firms. In contrast to conventional modes of disclosure, such as mandatory disclosures and traditional voluntary disclosures like conference calls, which follow specific procedures and formats, these novel online platforms enable direct, interactive, near real-time communication between firms and ordinary investors through the format of Q&A. By analyzing platform interactions, researchers can assess how effectively companies communicate with investors, what information they value, and how disclosures impact trading. This improves market regulation and corporate insights.
The interactive nature of these platforms provides a rich dataset to assess the quality of information disclosure. However, the sheer volume of data and the diversity in the quality of questions and responses pose significant challenges for large-scale analysis. Leuz and Wysocki Leuz and Wysocki (2016) summarize that the commonly used disclosure measures in the financial sector primarily capture the quantity rather than the quality of the information provided. However, it is possible to construct the quality dimension of disclosure by focusing on specific disclosures, and Natural Language Processing (NLP) emerges as a promising solution to evaluate the quality of information disclosure at scale. Existing applications of NLP in automatic financial and accounting text analysis have demonstrated success in various domains, such as automatic analysis of annual financial reports Smailović et al. (2017) and automatic sentiment analysis system on financial news articles Van de Kauter et al. (2015). However, to our knowledge, there is no standard dataset that evaluates the quality of financial information disclosure on these Q&A platforms.
In this work, we present a Q&A information disclosure quality assessment benchmark that evaluates Chinese Q&A data sourced from interactive platforms for communication between investors and listed companies, as established by SSE and SZSE. In developing this model, we have explored several technical challenges, including the interpretation of natural language, context sensitivity, and the subjective nature of relevance and readability assessments. This paper details the development, implementation, and evaluation of our NLP model, emphasizing its potential impact on the financial sector and the broader implications for automated quality assessment in investor communications. Datasets and code are available at https://anonymous.4open.science/r/FinTruthQA-218F.
2 Related work
Disclosure measurement
The measurement issue pertaining to disclosures has been a persistent challenge in empirical accounting research Leuz and Wysocki (2016). Predominantly, comprehensive measures have focused on quantifying disclosures, encompassing criteria such as the presence or absence of quarterly and annual reports, as well as numerical counts of activities like conference calls and earnings forecasts Lee and Zhong (2022). Another commonly used measure is the Association for Investment Management and Research (AIMR) rankings, based on annual surveys of financial analysts who evaluate U.S. firms’ disclosure practices Nagar et al. (2003); Healy et al. (1999); Welker (1995). However, they are limited to large U.S. firms and cover only a specific time period. Furthermore, alternative measures designed for specific disclosures typically employ a keyword-based approach to analyze the tone and sentiment of well-established, traditional disclosures such as earnings forecasts and conference call transcripts Bochkay et al. (2023); Frankel et al. (2010); Li (2010).
Financial NLP
Financial documents, such as annual reports, financial news articles, and economic policy statements, are typically composed in technical language in a non-structured text format and possess unique features distinct from other domains. Chang et al. (2016) Previous studies have focused on developing pre-trained models tailored for the financial domain Araci (2019); Yang et al. (2020); Liu et al. (2021). These efforts are predominantly oriented towards sentiment classification tasks in the financial context Wan et al. (2021). Additional research in this field has concentrated on question-answering tasks, involving datasets that combine textual and tabular information, where numerical analysis is often essential to derive answers Zhu et al. (2021); Zhao et al. (2022); Li et al. (2022). Recently, a series of financial LLMs has been developed for domain-specific applications Xie et al. (2024b); Bhatia et al. (2024). Xie et al. Xie et al. (2024a) introduced a comprehensive benchmark to evaluate the performance of LLMs across diverse financial tasks, and Aguda et al. Aguda et al. (2024) explored the use of LLMs as financial data annotators, with a primary emphasis on the English language. To our knowledge, there has been no prior research or dataset dedicated to the automatic quality assessment of financial Q&A data.
3 Dataset construction
3.1 Data collection
The data for this study was obtained from the interactive platforms for communication between investors and listed companies established by the Shanghai Stock Exchange (Shangzhen E-Hudong https://sns.sseinfo.com/) and the Shenzhen Stock Exchange (Hudongyi https://irm.cninfo.com.cn/). These two stock exchange sites cover almost all listed companies in China. These Q&A interactions are authentic, generated by human users and company representatives, and are regulated by the China Securities Regulatory Commission, ensuring the accuracy of disclosures. The companies involved are legally obligated to ensure the accuracy of their disclosures, and failure to do so could result in legal consequences. This regulatory oversight ensures that the data we used reflects genuine interactions. We collected over 900,000 Q&A entries, among which we randomly selected 6,000 samples for manual annotation, with 3,000 entries from each exchange’s platform. The data collection involved using web scraping techniques to extract the Q&A entries from the official platforms. The data were then manually reviewed and processed to ensure quality and consistency. The process was overseen by a PhD student from the Business School of XXX University. The ethics statement, including considerations of general ethical conduct and potential negative societal impacts, can be found in Supplementary Materials Section A. We also give several examples in Supplementary Materials Section G to demonstrate the wide range of investor inquiries covered in our dataset.
3.2 Data annotation
We focus on four key information disclosure quality evaluation criteria111In this paper, the terms ”criteria” and ”tasks” are used interchangeably. ”Criteria” refers to the specific standards used to evaluate the quality of the Q&A pairs, while ”tasks” refers to the individual evaluation processes performed based on these criteria, particularly in the experimental sections.: question identification, question relevance, answer relevance, and answer readability. These criteria are widely recognized as crucial indicators of information quality and are important for investors to consider when evaluating Q&A information. Specifically, the criteria in FinTruthQA are defined as follows:
1.
Question identification: This criterion assesses whether the question from an investor is a genuine question, as opposed to a suggestion or complaint. Each question is assigned a binary label: Positive, indicating it is a real question, or Negative, indicating it is not. Data entries with questions labeled as Negative are excluded from further evaluation under the remaining criteria.
2.
Question relevance: This criterion evaluates whether a question is pertinent to the investment topic, labeling it as Positive if relevant and Negative if not. It assesses whether the question is related to the company’s shareholders, financial indicators, industry conditions, or other investment-related topics.
3.
Answer readability: This criterion evaluates how easily an answer can be understood and whether it is well-written. It considers multiple dimensions, including structural coherence, clarity of information, stylistic appropriateness, and grammatical correctness, ensuring a comprehensive evaluation of the answer’s quality. The readability score ranges from 1 to 4, with 4 indicating the highest level of readability—clear, concise, and free of ambiguity. A score of 3 indicates moderate readability with some issues, while a score of 2 reflects significant readability challenges. A score of 1 suggests the answer is unreadable, with major ambiguities, incoherent sentences, or content unrelated to the question.
4.
Answer relevance: This criterion evaluates whether an answer is relevant and informative in response to the question. It assesses whether the answer provides sufficient detail and explanation to address the question. Answers are classified into three levels: Level 1 indicates the answer is completely unrelated to the question (none of the questions is answered); Level 2 indicates the answer is partially related (some questions are answered or it is difficult to judge) ; and Level 3 indicates the answer fully responds to the question (all questions are answered).
The importance of these criteria lies in their collective ability to enhance the quality and reliability of information on investor communication platforms. The annotation process was carried out in three phases: pre-annotation, formal annotation, and data verification and integration. In the pre-annotation phase, we recruited six students from the Accounting Department at XXX University and conducted five rounds of annotation. After the last round, the inter-annotator agreement (IAA) rate was calculated, and the four students with the highest average IAA rates (85.52%)—two undergraduates and two master’s students—were selected for the formal annotation phase. To ensure annotation quality, the four selected annotators were evenly divided into two groups, with each group annotating the same data instances. Discrepancies were first discussed and resolved within the group, with unresolved issues escalated to the project leader for final decisions. After the formal annotation, the project leader verified the data to ensure alignment with the latest annotation criteria and completeness, before integrating the data and summarising key information. Figure 2 illustrates an example of the annotation. A detailed description of the annotation process is listed in Supplementary Materials Section B.
3.3 Quality control
For the answer relevance task, we observed that the raw annotations contained noise due to the subjective nature of this task. To address this, we introduced an additional quality control step to correct the raw annotations. We implemented an enhanced correction process inspired by Confident Learning Northcutt et al. (2021) principles. Sample data points demonstrating inaccuracies, along with implementation details, can be found in Supplementary Materials Section C.
3.4 Dataset statistics
The complete annotated dataset comprises 6,000 samples labeled for question identification, question relevance, answer relevance, and answer readability tasks, with each data point having four distinct labels corresponding to these tasks. The dataset was randomly divided into training, validation, and testing sets in a ratio of 4:1:1. The label distributions for each task are detailed in Table 1. The provided statistics reflect the data after undergoing the quality control process. Note that our later analysis focuses exclusively on real questions. Only for the first task (question identification), we used the full set of 6,000 samples. For the remaining three tasks, we removed non-question samples, as they could skew the model’s understanding and performance, especially for tasks like question relevance, where the context of a real question is crucial for accurate assessment. It is worth noticing that the question relevance labels are highly imbalanced, with only 6 and 7 negative labels in the validation and test sets, respectively. We retained this task as the imbalance reflects the real-world data distribution, and the concept of question relevance is important in financial information disclosure. Figure 3 shows the length distribution of questions and answers, measured in characters, for all 6,000 QA pairs.
| Task | Label | # Train | # Valid | # Test | # Label | Freq. | Total |
| Question identification | Negative | 327 | 90 | 81 | 498 | 8.30% | 6,000 |
| Positive | 3,673 | 910 | 919 | 5,502 | 91.70% | ||
| Question relevance | Negative | 714 | 6 | 7 | 27 | 0.49% | 5,502 |
| Positive | 3,654 | 911 | 910 | 5,475 | 99.51% | ||
| Answer readability | 1 | 106 | 190 | 198 | 1,094 | 19.88% | 5,502 |
| 2 | 26 | 3 | 6 | 35 | 0.64% | ||
| 3 | 68 | 19 | 20 | 107 | 1.94% | ||
| 4 | 2,868 | 705 | 693 | 4,266 | 77.54% | ||
| Answer relevance | 1 | 731 | 194 | 206 | 1,131 | 20.63% | 5502 |
| 2 | 533 | 136 | 130 | 779 | 14.21% | ||
| 3 | 2,404 | 587 | 581 | 3,572 | 65.16% |
4 Benchmarks
4.1 Model selection
To fully evaluate the performance of different methods in FinTruthQA, we benchmarked several models that cover Machine Learning (ML)-based, Pre-trained Language Model (PLM)-based, PLM with continued Pre-Training (CPT)-based, and LLM-based approaches. Given the distinct nature and varying difficulty of the tasks, we applied tailored approaches to each accordingly.
ML-based. We modeled these tasks as text classification problems and used various advanced machine learning models, including Logistic Regression (LR) Hosmer Jr et al. (2013), Support Vector Machine (SVM) Hearst et al. (1998), K-nearest neighbor (KNN) Peterson (2009), and Random Forest (RF) Breiman (2001). For semantic feature extraction, we first utilized the Jieba toolkit 222https://github.com/fxsjy/jieba for Chinese word segmentation and then extracted weighted text representations using TF-IDF Sparck Jones (1972). To achieve optimal performance, we used cross-validation and GridSearch to rigorously explore the best hyperparameters. The best-performing classifiers and hyperparameters varied across tasks. Therefore, we performed task-specific hyperparameter searches for each task individually.
PLM-based. In our experiments, we found that the tasks of question identification and question relevance were relatively straightforward, with BERT achieving excellent performance on these tasks. Therefore, our focus shifted to exploring the effectiveness of PLM-based methods on the more challenging tasks of answer relevance and answer readability. Specifically, we investigated a range of BERT-based models, including those specifically trained on financial text data, such as financial news, research reports, and financial encyclopedia entries (FinBERT,333https://github.com/valuesimplex/FinBERT Mengzi-fin Zhang et al. (2021), and DKPLM Zhang et al. (2022)). Furthermore, we examined whether fine-tuned PLMs on specific tasks, such as NLI (Natural Language Inference) and QA (Question Answering), would offer any benefits for these tasks (SBERT-nli Reimers and Gurevych (2019) and RoBERTa-extractive-qa Liu et al. (2019)). Finally, we also examined the performance of a large BERT model (BERT (Large)).444https://huggingface.co/yechen/bert-large-chinese Notably, unless otherwise specified, all PLMs employed in this study are the base and Chinese versions. Links for these PLMs can be found in Supplementary Materials Section D.
PLM with CPT-based. General domain PLMs, trained on broad texts, often lack the specialized knowledge required to handle the nuanced language and terminology in finance. This limitation is especially pronounced in tasks like answer relevance, where understanding domain-specific jargon and abbreviations is essential for accurate predictions. To bridge this gap, we curated a substantial volume of domain-specific texts and conducted continued pre-training on general PLMs. Specifically, we collected over 900,000 Q&A pairs from the SSE and SZSE platforms by web scraping. The top-performing models from previous experiments were then selected for continued pre-training using masked language modeling (MLM). This pre-training on a large-scale, domain-specific dataset enhances the PLM’s ability to understand financial contexts, capture more specialized information, and achieve better performance in downstream tasks.
LLM-based. We further evaluate some advanced LLMs, including GPT-4 Achiam et al. (2023), LLaMA-3.1 Dubey et al. (2024), Qwen2 Yang et al. (2024), and Mistral Jiang et al. (2023), on the most challenging answer relevance task. Due to their in-context learning ability, LLMs can perform this task in a zero-shot manner, without requiring specific fine-tuning. Specifically, we provided LLMs with task descriptions and evaluation criteria, instructing them to directly output relevance scores (Direct score). Moreover, we incorporated the Chain of Thought (CoT) Wei et al. (2022) technique, prompting the LLMs to provide specific reasoning alongside the relevance scores (Score with CoT). Full details are provided in Supplementary Section F.
4.2 Settings and metrics
Implementation details. For the ML-based models, we utilized the standard scikit-learn555https://scikit-learn.org/stable/ library. For all PLM-based methods, we used PyTorch666https://github.com/pytorch/pytorch and HuggingFace’s framework777https://github.com/HuggingFace for downstream fine-tuning and evaluation. Detailed hyperparameter settings for our PLM-based experiments can be found in Supplementary Materials Section E. We conducted experiments comparing Q&A pairs with and without the company name as a prefix and found negligible differences, so we ultimately used data without the prefix. The tasks of question identification and question relevance only required encoding the questions as input, while answer readability and answer relevance required encoding both the question and the answer. The sequence started with a special classification token [CLS], and the question and answer were separated by the [SEP] token . During continued pre-training, we added two special tokens, [Q] and [A], into the vocabulary and prefixed them to the question and answer, respectively. This helps the PLMs better identify the components of each sentence. The continued pre-training was performed based on the UER framework Zhao et al. (2019). All tests of LLMs were conducted using the transformer framework888https://github.com/huggingface/transformers, except for the GPT-4 experiments, which were conducted by calling OpenAI’s API with version GPT-4-0613. Specifically, we used Llama-3.1 with 8 billion parameters, and both Qwen2 and Mistral with 7 billion parameters.
| Category | Model | Question identification (%) | Question relevance(%) | ||||||
| Accuracy | Precision | Recall | F1 | Accuracy | Precision | Recall | F1 | ||
| ML-based | TF-IDF + LR | 93.700.86 | 94.110.75 | 99.350.18 | 96.660.48 | 99.420.13 | 99.420.13 | 100.000.00 | 99.710.07 |
| TF-IDF + RF | 93.700.65 | 94.020.63 | 99.460.18 | 96.66 0.37 | 99.420.13 | 99.420.13 | 99.930.10 | 99.670.08 | |
| TF-IDF + SVM | 93.600.73 | 93.780.76 | 99.640.13 | 96.62 0.40 | 99.420.13 | 99.420.13 | 100.000.00 | 99.710.07 | |
| TF-IDF + KNN | 92.671.11 | 93.270.98 | 99.160.37 | 96.120.61 | 99.420.13 | 99.420.13 | 100.000.00 | 99.710.07 | |
| PLM-based | BERT | 96.890.55 | 97.640.22 | 99.020.41 | 98.320.28 | 99.670.11 | 99.720.12 | 99.960.05 | 99.850.06 |
Metrics
For question identification and question relevance, which are binary classification tasks, we used accuracy, precision, recall, and F1-score as evaluation metrics. For answer readability and answer relevance, which are multi-class classification tasks, we calculated both micro and macro F1-scores, along with the Quadratic Weighted Kappa (QWK) Dong et al. (2017), which is well-suited for ordinal classification by penalizing misclassifications based on their distance from the true label. Each model was run in five times, and the mean and variance of the results are reported.
4.3 Results and analysis
4.3.1 Task 1: Question identification
| Category | Model | F1 (Micro) | F1 (Macro) | F1 (Weighted) | QWK |
| ML-based | TF-IDF + LR | 81.72 0.27 | 34.910.22 | 78.900.27 | 41.200.80 |
| TF-IDF + RandomForest | 80.480.70 | 32.811.06 | 76.851.03 | 33.793.87 | |
| TF-IDF + SVM | 81.640.37 | 33.920.38 | 78.15 0.40 | 37.981.48 | |
| TF-IDF + KNN | 78.990.34 | 31.071.23 | 74.95 0.84 | 27.314.28 | |
| PLM-based | BERT | 86.040.44 | 41.970.67 | 84.890.44 | 63.521.43 |
| RoBERTa | 86.260.62 | 42.021.92 | 85.080.89 | 64.000.96 | |
| FinBERT | 87.570.39 | 41.340.45 | 85.980.69 | 66.531.83 | |
| Mengzi-fin | 86.990.80 | 41.931.42 | 85.771.20 | 66.331.81 | |
| DKPLM | 87.240.76 | 41.901.17 | 85.760.72 | 65.780.63 | |
| SBERT-nli | 85.311.05 | 39.280.81 | 83.481.38 | 58.192.77 | |
| RoBERTa-extractive-qa | 87.210.52 | 41.220.11 | 85.750.64 | 65.970.80 | |
| BERT (Large) | 87.241.03 | 42.171.66 | 85.781.30 | 65.922.74 | |
| RoBERTa (Large) | 86.481.48 | 45.514.59 | 85.371.41 | 64.582.11 |
In this task, we benchmarked several ML-based and PLM-based models, with their performances summarized in Table 2. The ML-based models showed consistent results with minor variations. Both TF-IDF + LR and TF-IDF + RF reached an accuracy of 93.70% and an F1 score of 96.66, with closely aligned precision and recall. TF-IDF + SVM exhibited slightly lower accuracy but achieved the highest recall among the ML models, demonstrating its strength in identifying relevant instances. TF-IDF + KNN, while performing well with an accuracy of 92.67%, had the lowest precision and F1 score in this group. On the other hand, the PLM-based BERT model outperformed all ML-based methods, achieving the highest accuracy, precision, and F1 score. This result highlights the PLM model’s ability to capture deeper contextual nuances, which traditional frequency-based methods like TF-IDF struggle to achieve.
4.3.2 Task 2: Question relevance
Table 2 presents the evaluation results for the question relevance task. The PLM-based model outperformed the ML-based models on all metrics except recall, achieving an accuracy of 99.67%, an F1 score of 99.85%, precision of 99.72%, and recall of 99.96%. This demonstrates its ability to effectively identify relevant questions while maintaining a strong balance in correctly classifying rare negative samples. In contrast, the TF-IDF-based models showed nearly identical performance across metrics, primarily predicting all samples as positive. Given the highly imbalanced label distribution (910 positive vs. 7 negative samples in the test set), these models achieve high recall by predicting positives but struggle to correctly identify the few negative instances, reflecting the challenge posed by the skewed data.
4.3.3 Task 3: Answer readability
The evaluation results for the task answer readability are shown in Table 3. Among the ML-based models, the LR classifier exhibited the highest performance, achieving an F1 Micro score of 81.72% and a QWK of 41.20%. While LR provided reasonable performance, the relatively low F1 Macro score highlights its difficulties in handling class imbalance effectively. In contrast, the PLM-based models demonstrated significant improvements over the ML-based models. For instance, BERT achieved an F1 Micro score of 86.04% and a QWK of 63.52%, showcasing its superior ability to capture the complexities of readability assessment through richer contextual understanding.
Detailed Evaluation of FinBERT’s Performance.
FinBERT outperformed all other models with an F1 Micro score of 87.57%, an F1 Macro score of 41.34%, an F1 Weighted score of 85.98%, and the highest QWK of 66.53%. Its pre-training on financial text provided a distinct advantage for this domain-specific task, enabling it to achieve the highest accuracy and alignment with human judgments. In further analysis, we find that the model performs well in recognizing answers with the highest readability, but it struggles to accurately differentiate between the lower readability levels, highlighting the need for improvement in capturing finer distinctions in readability.
4.3.4 Task 4: Answer relevance
| Category | Model | F1 (Micro) | F1 (Macro) | F1 (Weighted) | QWK |
| ML-based | TF-IDF + LR | 69.610.49 | 49.740.53 | 65.070.60 | 38.162.13 |
| TF-IDF + RF | 69.970.34 | 44.79 0.84 | 63.070.76 | 36.631.10 | |
| TF-IDF + SVM | 70.080.83 | 44.420.43 | 62.931.07 | 36.781.65 | |
| TF-IDF + KNN | 66.661.03 | 38.632.05 | 57.871.42 | 23.223.36 | |
| PLM-based | BERT | 77.900.98 | 67.640.94 | 76.790.75 | 63.950.85 |
| RoBERTa | 77.101.34 | 65.952.64 | 75.771.35 | 62.222.46 | |
| FinBERT | 78.190.96 | 68.180.79 | 77.250.81 | 65.490.54 | |
| Mengzi-fin | 78.040.72 | 68.962.42 | 77.440.82 | 66.311.10 | |
| DKPLM | 76.920.57 | 63.00 3.72 | 74.651.15 | 62.821.78 | |
| SBERT-nli | 74.520.36 | 61.182.54 | 72.381.07 | 53.881.12 | |
| RoBERTa-extractive-qa | 78.330.36 | 67.021.48 | 77.120.69 | 64.661.44 | |
| BERT (Large) | 78.010.78 | 65.941.62 | 76.160.56 | 63.530.90 | |
| RoBERTa (Large) | 78.300.94 | 68.502.09 | 77.401.54 | 64.521.56 | |
| PLM with CPT-based | FinBERT | 80.040.23 | 70.580.90 | 79.020.18 | 67.891.26 |
| RoBERTa-extractive-qa | 78.340.34 | 68.01 1.71 | 77.450.74 | 66.760.58 | |
| BERT (Large) | 78.12 0.65 | 65.153.97 | 76.02 1.16 | 64.822.07 | |
| RoBERTa (Large) | 74.371.78 | 61.401.28 | 72.460.73 | 53.913.42 | |
| LLM-based | GPT-4 (Direct score) | 67.39 | 63.11 | 70.14 | 59.06 |
| GPT-4 (with CoT) | 59.76 | 58.45 | 64.28 | 57.51 | |
| Llama-3.1 (Direct score) | 53.00 | 41.59 | 56.13 | 43.24 | |
| Llama-3.1 (with CoT) | 46.46 | 40.82 | 51.57 | 39.56 | |
| Qwen2 (Direct score) | 60.52 | 31.33 | 52.83 | 23.68 | |
| Qwen2 (with CoT) | 59.00 | 38.97 | 57.36 | 42.35 | |
| Mistral (Direct score) | 54.53 | 30.40 | 50.99 | 29.57 | |
| Mistral (with CoT) | 59.32 | 23.03 | 60.20 | 38.90 |
PLM-based Methods Consistently Outperform ML-based ones.
As shown in Table 4, all PLM-based methods can achieve better performance than ML-based methods across all metrics. For instance, RoBERTa-extractive-qa achieved the highest Micro F1 score of 78.33%, while Mengzi-fin excelled with the highest Macro F1 score of 68.96% and a strong QWK of 66.31%. Their superior performance can be attributed to pre-training on task-specific and domain-specific data: RoBERTa-extractive-qa was pre-trained on QA datasets, which enhanced its ability to comprehend and evaluate answer relevance, while Mengzi-fin was pre-trained on financial corpora, enabling it to better capture domain-specific nuances in the financial context.
Impact of Continued Pre-training on PLM-based Methods.
We selected four PLM-based models and further pre-trained them using 900,000 Q&A pairs. As shown in Table 4, we observed that only FinBERT demonstrated significant performance improvements, achieving the highest Micro F1 score of 80.04% and a QWK of 67.89%, indicating its enhanced ability to assess answer relevance in financial contexts. In contrast, the other three models (RoBERTa-extractive-qa, BERT Large, and RoBERTa Large) did not exhibit consistent improvements. This suggests that domain alignment plays a crucial role in the effectiveness of continued pre-training, as FinBERT’s domain-specific pre-training on financial texts provided a strong foundation, whereas the other models may have experienced a mismatch between their learned representations and the specific requirements of the task, resulting in less notable gains.
Suboptimal Performance of LLMs in Chinese Financial Contexts.
UTF8gbsn As shown in Table 4, we also explored the performance of advanced LLMs on answer relevance task. GPT-4 (Direct score) consistently outperforms other LLMs but still lags behind supervised PLM-based models, underscoring the difficulty LLMs face in understanding Chinese financial disclosures. Notably, GPT-4 with CoT did not outperform the GPT-4 Direct score. A potential reason is that GPT-4 was primarily trained on English corpora; it lacks sufficient financial domain knowledge in Chinese. This may result in knowledge gaps and hallucinations during the CoT process, thereby reducing the model’s performance. Another study also observed a similar phenomenon where CoT failed in healthcare domain prediction with Chinese text. Liu et al. (2024) We also tested Chinese financial LLMs, including CFGPT2 Li et al. (2023) and DISC-FinLLM Chen et al. (2023), but found that these models could hardly follow instructions and failed to generate reasonable predictions. Outputs often exhibited excessive repetition and encoding errors. As a result, we decided to exclude them from our study. We anticipate that further domain adaptation methods, such as few-shot learning or retrieval-augmented generation Lewis et al. (2020), may improve the performance of LLMs on this task. Figure 4(a) shows the confusion matrix for the best-performing model (FinBERT with continued pretraining). It demonstrates that the model effectively captures answers that are completely related to the questions (labeled as 3). However, the model struggles to distinguish partially unrelated answers (labeled as 2) from completely related ones, where the nuanced differences are subtle and challenging. The confusion matrices for GPT-4 (direct score) are shown in Figures 4(b), which illustrate that the model tends to mix up answers that are completely irrelevant (labeled as 1) with those that are partially relevant (labeled as 2). Also, there’s a tendency for the model to predict label 2. We also conducted a thorough examination of bad cases to better understand the prevalent error types exhibited by GPT-4 in relevance scoring, which can be found in Supplementary Section H.
5 Limitations
There are several limitations in this study. First, the Chinese cultural and regulatory context of SSE and SZSE may limit the generalizability of the findings to other markets. Future research is needed to explore the applicability of the developed models and methodologies in different contexts and validate the findings across diverse settings. Second, the annotation process inherently involves a degree of subjectivity, as differences in annotators’ interpretations can introduce variability in the labeled data. To mitigate this, we conducted multiple rounds of pre-annotation and formal annotation with professional annotators, along with thorough error analysis and expert reviews to ensure consistency. Third, while precision, recall, F1-score, and quadratic weighted kappa (QWK) are standard metrics for evaluating model performance, they may not fully capture the qualitative aspects of information disclosure quality. The reliance on these metrics could lead to an overemphasis on quantitative evaluation while neglecting interpretability. Third, given the substantial computational cost of training, we limited LLM evaluations to a zero-shot setting for answer relevance. This decision balances feasibility with the need to showcase LLM capabilities. Future studies will incorporate fine-tuned evaluations of LLMs to provide a more consistent comparative framework. Lastly, the question relevance task suffers from a highly imbalanced dataset, which may limit the model’s ability to effectively learn from and evaluate rare instances. However, we chose to retain this task in the benchmark to reflect the real-world data distribution in financial information disclosure.
6 Conclusion
This study introduces FinTruthQA, a benchmark designed to evaluate advanced NLP techniques for the automatic quality assessment of information disclosure in financial Q&A data. We constructed a dataset of 6,000 real-world financial Q&A entries, each manually annotated based on four key evaluation criteria of accounting. Experiments revealed the effectiveness of pre-trained BERT models in tackling these tasks. In particular, we observed that models pre-trained on financial corpora, as well as those designed for related tasks such as QA, outperformed other variants of BERT and GPT-4. This emphasized the importance of domain-specific and task-oriented training in improving performance. Although existing NLP models showed reasonable performance on this benchmark, there is still significant room for improvement, especially in the task of answer relevance, which calls for better NLP models. Our benchmark is not only useful for post-hoc regulatory assessments but also holds significant value in real-time applications, highlighting the societal impact of AI in fostering financial transparency, reducing information asymmetry, and promoting fairer market participation. For instance, auditors, financial analysts, and corporate communication teams can leverage our framework for ongoing monitoring and assessment of information disclosure quality. By facilitating real-time identification and correction of inadequate disclosures, this application enhances market efficiency, promotes transparency, and fosters equitable access to financial information. Ultimately, it supports fairer investment opportunities, strengthens investor confidence, and contributes to sustainable economic development.